xx NullExtract

NullExtract

Maria Babiuc-Hamilton <babiuc@marshall.edu>

June 26 2009

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11afact 10Scope: privatex01 REAL

11Description: Factor in front of the sine time dependence for Schwazchild time metric

11 Range 10 x01 Default: 0.0
11*:*  10anything

11

11anaxmetric  10Scope: privatex01  BOOLEAN

11Description: Should the Schwarzchild analytic metric be used?

11                           10  x01  Default: no

11

11debugxswitchxon  10Scope: privatex01  BOOLEAN

11Description: Should the debug be turned on?

11                           10  x01  Default: no

11

11elldxminxcoef  10Scope: privatex01  REAL

11Description: definition of elldxmin = elldxminxcoef * dt**elldxminxpow

11 Range 10 x01 Default: 0.01
110:*  10larger than zero

11

11elldxminxpow  10Scope: privatex01  INT

11Description: definition of elldxmin = elldxminxcoef * dt**elldxminxpow

11 Range 10 x01 Default: 2
110:*  10larger than zero

11

11fcoef  10Scope: privatex01  REAL

11Description: Coefficient of the time dependence for Schwazchild time metric

11 Range 10 x01 Default: 0.0
11*:*  10anything

11

11haltxonxnegativexelld  10Scope: privatex01  BOOLEAN

11Description: Should the code quit if the denominator of the null vector changes sign?

11                           10  x01  Default: no

11

11haltxonxnegativexrl  10Scope: privatex01  BOOLEAN

11Description: Should the code quit if rx{lambda} goes negative?

11                           10  x01  Default: no

11

11rlxminxcoef  10Scope: privatex01  REAL

11Description: definition of rlxmin = rlxminxcoef * dt**rlxminxpow

11 Range 10 x01 Default: 0.01
110:*  10larger than zero

11

11rlxminxpow  10Scope: privatex01  INT

11Description: definition of rlxmin = rlxminxcoef * dt**rlxminxpow

11 Range 10 x01 Default: 2
110:*  10larger than zero

11schiefxtime  10Scope: privatex01  KEYWORD

11Description: Should the analytic metric be static or sin(omega t)

11 Range 10 x01 Default: static
11static  10Static solution, no time dependence
11sinext  10Sine dependence in time
11vibext  10A sine/cosine vibe in time on the x direction

11

11wtxmetric  10Scope: privatex01  KEYWORD

11Description: Metric on the Worldtube to be used

11 Range 10 x01 Default: Full
11Full  10Full nonlinear metric extraction on the worldtube
11SchIEF  10Schwarzchild metric in IEF coordinates

11

11wtxsphericalxharmonics  10Scope: privatex01  BOOLEAN

11Description: Output the spherical harmonics of the Worldtube to a file?

11                           10  x01  Default: no

11

11wtxsphericalxharmonicsxonxnullgrid 10Scope: privatex01  BOOLEAN

11Description: Output the spherical harmonics on the nullgrid just off the Worldtube to a file?

11                           10  x01  Default: no

11

11cr  10Scope: restrictedx01  REAL

11Description: radius of the worltube

11 Range 10 x01 Default: 0.5
11(0.0:*  10strictly positive real number

11

11lxmax  10Scope: restrictedx01  INT

11Description: lxmax in Ylm

11 Range 10 x01 Default: 2
110:*  10positive

11

11mass  10Scope: restrictedx01  REAL

11Description: mass parameter

11 Range 10 x01 Default: 0.1
11(0.0:*  10strictly positive real number

11

11minimumdistanceto2bp1  10Scope: restrictedx01  REAL

11Description: minimum distance from the extraction world-tube to the B+1 point

11 Range 10 x01 Default: 0.0
110:*  10positive. Should be O(1)

11

11nxangxevxoutsidexeq  10Scope: shared from NULLGRIDx01INT

11

11nxangxghostxpts  10Scope: shared from NULLGRIDx01INT

11

11nxangxptsxinsidexeq  10Scope: shared from NULLGRIDx01INT

11

11nxangxstencilxsize  10Scope: shared from NULLGRIDx01INT

11nxradialxpts  10Scope: shared from NULLGRIDx01INT

11

11nullxrwt  10Scope: shared from NULLGRIDx01REAL

11

11nullxxin  10Scope: shared from NULLGRIDx01REAL

11

7 Interfaces

General Implements: nullshrextract

Inherits: nullinterp nullgrid nullvars

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  nylm  10  x01  compact110
11                 10  reYN  x01 description11Ylm in (p
11                 10  x  x01 description11q) coordinates – north patch
11                 10  imYN  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  sylm  10  x01  compact110
11                 10  reYS  x01 description11Ylm in (p
11                 10  x  x01 description11q) coordinates – south patch
11                 10  imYS  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  n1ylm  10  x01  compact110
11                 10  re1YN  x01 description111Ylm in (p
11                 10  x  x01 description11q) coordinates – north patch
11                 10  im1YN  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  s1ylm  10  x01  compact110
11                 10  re1YS  x01 description111Ylm in (p
11                 10  x  x01 description11q) coordinates – south patch
11                 10  im1YS  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  ntylm  10  x01  compact110
11                 10  reTYN  x01 description11Ylm in (p
11                 10  x  x01 description11q) coordinates for Schw metric north patch
11                 10  imTYN  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  stylm  10  x01  compact110
11                 10  reTYS  x01 description11Ylm in (p
11                 10  x  x01 description11q) coordinates for Schw metric south patch
11                 10  imTYS  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11  nt1ylm  10  x01  compact110
11                 10  re1TYN  x01 description111Ylm in (p
11                 10  x  x01 description11q) coordinates for Schw metric north patch
11                 10  im1TYN  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  st1ylm  10  x01  compact110
11                 10  re1TYS  x01 description111Ylm in (p
11                 10  x  x01 description11q) coordinates for Schw metric south patch
11                 10  im1TYS  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  tclm  10  x01  compact110
11                 10  RTC  x01 description11Expansion coefficients for the Schwarzchild metric
11                 10  ITC  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  tcrlm  10  x01  compact110
11                 10  RTCr  x01 description11Radial derivatives of expansion coefficients for the Schwarzchild metric
11                 10  ITCr  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  tctlm  10  x01  compact110
11                 10  RTCt  x01 description11Time derivatives of xpansion coefficients for the Schwarzchild metric
11                 10  ITCt  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01vararrayxsize11(lxmax+1)*(lxmax+1)
11                 10  x01variable type11REAL
11  clm  10  x01  compact110
11                 10  RC  x01 description11Expansion coefficients for the metric
11                 10  IC  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11  drclm  10  x01  compact110
11                 10  RCr  x01 description11Expansion coefficients for the radial derivative of the metric
11                 10  ICr  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11  dtclm  10  x01  compact110
11                 10  RCt  x01 description11Expansion coefficients for the time derivative of the metric
11                 10  ICt  x01 dimensions111
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11ARRAY

11                 10  x01  size1110

11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 wtxr1 10 WTxr1 x01 compact110
11                 10  x01 description11Lambda derivative of the radius
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11  wtxscalars  10  x01  compact110
11                 10  WTxdetg  x01 description11various scalars
11                 10  WTxtemp  x01 dimensions112
11                 10  WTxsigma2  x01 distribution11DEFAULT
11                 10  WTxelld  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 wtx3vectors 10 x01 compact110
11                 10 WTxsigmaxvect x01 description11various 3-vectors
11                 10 WTxdr0xvect x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize116
11                 10  x01variable type11REAL
11 wtx4vectors 10 x01 compact110
11                 10 WTxdr1xvect x01 description11various 4-vectors
11 10 WTxellxvect x01 dimensions112
11                 10 WTxdelldxvect x01 distribution11DEFAULT
11                 10  WTxnaxvect  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  WTxsaxvect  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize118
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 wtxd4vectors 10 x01 compact110
11                 10 WTxDSnormxvect x01 description113-deriv of various 4-vectors
11                 10 WTxDNnormxvect x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1124
11                 10  x01variable type11REAL
11 wtxdsigmaxvect 10 WTxdsigmaxvect x01  compact110
11                 10  x01 description11derivative of sigma
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11 wtxg1xvect 10 WTxg1xvect x01 compact110
11                 10  x01 description11lambda deriv. of covariant metric
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1120
11                 10  x01variable type11REAL
11 wtxgupxvect 10 WTxgupxvect x01  compact110
11                 10  x01 description11contravariant metric
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1120
11                 10  x01variable type11REAL
11  wtxjac  10  x01  compact110
11                 10  WTxj0xvect  x01 description11jacobian
11                 10 WTxj0invxvect x01 dimensions112
11 10 WTxj1xvect x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1132
11                 10  x01variable type11REAL
11 wtxdj0xvect 10 WTxdj0xvect x01  compact110
11                 10  x01 description11second angular deriv. of Cart. coord
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1118
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 wtxaffinexmetric 10 x01 compact110
11                 10 WTxeta0xvect x01 description11affine metric and its derivative
11                 10 WTxetaup0xvect x01 dimensions112
11                 10 WTxeta1xvect x01 distribution11DEFAULT
11                 10 WTxetaup1xvect x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1120
11                 10  x01variable type11REAL
11 wtxdeta0xvect 10 WTxdeta0xvect x01  compact110
11                 10  x01 description11Angular derivatives of angular null metric components
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11 wtxcmpxcoord 10  x01  compact110
11                 10  qax2  x01 description11The complex dyad
11                 10  qax3  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01variable type11COMPLEX

11

7.0.2 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 shrexgij 10 SHRExgij x01 compact110
11                 10  x01 description11The spatial metric components on the extraction worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11 shrexgit 10 SHRExgit x01 compact110
11                 10  x01 description11The mixt space-time metric components on the extraction worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize118
11                 10  x01variable type11REAL
11 shrexbeta 10 SHRExbeta x01 compact110
11                 10  x01 description11The shift beta
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize116
11                 10  x01variable type11REAL
11 shrexalpha 10 SHRExalpha x01 compact110
11                 10  x01 description11The lapse alpha
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11  shrexdgij  10  x01  compact110
11                 10  SHRExdrgij  x01 description11Derivatives of the spatial metric components on the extraction worldtube
11                 10  SHRExdqgij  x01 dimensions112
11                 10  SHRExdpgij  x01 distribution11DEFAULT
11                 10  SHRExdtgij  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11  shrexdgit  10  x01  compact110
11                 10  SHRExdrgit  x01 description11Derivatives of the mixt metric components on the extraction worldtube
11                 10  SHRExdqgit  x01 dimensions112
11                 10  SHRExdpgit  x01 distribution11DEFAULT
11                 10  SHRExdtgit  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize118
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 shrexdbeta 10 x01 compact110
11                 10 SHRExdrbeta x01 description11Derivatives of the shift beta
11                 10 SHRExdqbeta x01 dimensions112
11                 10 SHRExdpbeta x01 distribution11DEFAULT
11                 10 SHRExdtbeta x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize116
11                 10  x01variable type11REAL
11 shrexdalpha 10 x01 compact110
11                 10 SHRExdralpha x01 description11Derivatives of the lapse alpha
11                 10 SHRExdqalpha x01 dimensions112
11                 10 SHRExdpalpha x01 distribution11DEFAULT
11                 10 SHRExdtalpha x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels113
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11

7.0.3 PROTECTED GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 wtxr0 10 WTxr0 x01 compact110
11                 10  x01 description11Bondi radius of the world-tube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 xxwt 10 xxwt x01 compact110
11                 10  x01 description11The worldline path
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 jxwt 10 jxwt x01 compact110
11                 10  x01 description11The metric function J of the worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 jxl 10 jxl x01 compact110
11                 10  x01 description11Lambda derivative of the metric function J of the worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 betaxwt 10 betaxwt x01 compact110
11                 10  x01 description11The expansion factor beta
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 betaxl 10 betaxl x01 compact110
11                 10  x01 description11Lambda derivative of the expansion factor beta
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 uxwt 10 uxwt x01 compact110
11                 10  x01 description11The shift U
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 uxl 10 uxl x01 compact110
11                 10  x01 description11Lambda derivative of the shift U
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 uxxxwt 10 uxxxwt x01 compact110
11                 10  x01 description11Radial derivative of the shift U
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 qxwt 10 qxwt x01 compact110
11                 10  x01 description11The auxiliary variable of the U equation
11 10 x x01 description11Q
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11COMPLEX
11 wxwt 10 wxwt x01 compact110
11                 10  x01 description11The mass aspect W
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels112
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 wxl 10 wxl x01 compact110
11                 10  x01 description11Lambda derivative of the mass aspect W
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 shrexgijxe 10 SHRExgijxe x01 compact110
11                 10  x01 description11The spatial metric components on the extraction worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11 shrexgitxe 10 SHRExgitxe x01 compact110
11                 10  x01 description11The mixt space-time metric components on the extraction worldtube
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize118
11                 10  x01variable type11REAL
11 shrexbetaxe 10 SHRExbetaxe x01  compact110
11                 10  x01 description11The shift beta
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize116
11                 10  x01variable type11REAL
11 shrexalphaxe 10 SHRExalphaxe x01  compact110
11                 10  x01 description11The lapse alpha
11                 10  x01 dimensions112
11                 10  x01 distribution11DEFAULT
11                 10  x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL
11 shrexdgijxe 10 x01 compact110
11                 10 SHRExdrgijxe x01 description11Derivatives of the spatial metric components on the extraction worldtube
11                 10 SHRExdqgijxe x01 dimensions112
11                 10 SHRExdpgijxe x01 distribution11DEFAULT
11                 10 SHRExdtgijxe x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1112
11                 10  x01variable type11REAL
11 shrexdgitxe 10 x01 compact110
11                 10 SHRExdrgitxe x01 description11Derivatives of the mixt metric components on the extraction worldtube
11                 10 SHRExdqgitxe x01 dimensions112
11                 10 SHRExdpgitxe x01 distribution11DEFAULT
11                 10 SHRExdtgitxe x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize118
11                 10  x01variable type11REAL

11x Group Names x10x Variable Names xx01 Details x11x
11 shrexdbetaxe 10 x01 compact110
11                 10 SHRExdrbetaxe x01 description11Derivatives of the shift beta
11                 10 SHRExdqbetaxe x01 dimensions112
11                 10 SHRExdpbetaxe x01 distribution11DEFAULT
11                 10 SHRExdtbetaxe x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize116
11                 10  x01variable type11REAL
11 shrexdalphaxe 10 x01 compact110
11                 10 SHRExdralphaxe x01 description11Derivatives of the lapse alpha
11                 10 SHRExdqalphaxe x01 dimensions112
11                 10 SHRExdpalphaxe x01 distribution11DEFAULT
11                 10 SHRExdtalphaxe x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11 10 x x01 ghostsize11NULLGRID::NxANGxGHOSTxPTS
11                 10  x01 group type11ARRAY
11                 10  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x  x01  size11(NULLGRID::NxANGxPTSxINSIDExEQ+2*(NULLGRID::NxANGxEVxOUTSIDExEQ+NULLGRID::NxANGxSTENCILxSIZE))
11                 10  x01 timelevels111
11                 10  x01vararrayxsize112
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn PITTNullCode/NullSHRExtract. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage

11x11Conditional:
11x 11SHRExgij[3] SHRExgit[3] SHRExbeta[3] SHRExalpha[3]
11x 11SHRExdgij[3] SHRExdgit[3] SHRExdbeta[3] SHRExdalpha[3]
11x 11WTxjac WTxaffinexmetric WTxdeta0xvect
11x 11xxwt[2] jxwt[2] jxl betaxwt[2] betaxl
11x 11qxwt[2] uxwt[2] uxl uxxxwt wxwt[2] wxl
11x 11NYlm SYlm N1Ylm S1Ylm
11x 11Clm drClm dtClm
11x 11NTYlm STYlm NT1Ylm ST1Ylm
11x 11TClm TCtlm TCrlm
11x 11WTxr0 WTxr1[3] WTxdj0xvect
11x 11NullVars::realcharfuncs[2]
11x 11NullVars::cmplxcharfuncsxbasic[2]
11x 11NullVars::nullxmask
11x 11NullVars::cmplxcharfuncsxaux[2]
11x 11WTxCmpxCoord
11x 11WTxscalars WTx3vectors WTx4vectors WTxd4vectors
11x 11WTxdsigmaxvect WTxg1xvect WTxgupxvect

11x 11x

11

Scheduled Functions

MetricReconstruction (conditional) x nullshrexfullmetricrecon x current extraction of full metric and its derivatives on the sphere x

11x11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxINITIAL (conditional) x nullshrexprecalcnsylm x precalculate spherical harmonics in (p,q) coordinates x

11x11Before: 11nullevolxboundaryinit
11 x 11x  11nullevolxinitial
11 x 11Language:11fortran
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexpointingxpxp x target p-p level variables for the pointers used in initialization x

11x11Before: 11nullshrexinitialwtxpxp
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexinitialwt x extracted variables on the world tube for the pxpxinit level only up to rl

11x11After: 11nullshrexpointingxpxp
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexpointingxp x target p-level variables for the pointers used in initialization x

11x11After: 11nullshrexinitialwtxpxp
11 x 11Before: 11nullshrexinitialwtxp
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexinitialwt x extracted variables on the world tube for the p-init level only up to rl x

11x11After: 11nullshrexpointingxp
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexpointing x gives the target variables for the pointers used in initialization x

11x11After: 11nullshrexinitialwtxp
11 x 11Before: 11nullshrexinitialwt
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexextractwt x computes the extracted variables on the world tube for the initial level x

11x11After: 11nullshrexpointing
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullEvolxBoundary x nullshrextract x group of the waveform extraction current level

11x11Type:11group

11

NullEvolxBoundary (conditional) x sphericalharmonicreconread x schedule group for reconstructing current world-tube data x

11x11Before:11nullshrextract
11 x 11Type: 11group

11

NullSHRExtract (conditional) x nullshrexschreadclm x current schwarzchild spherical harmonic coefficients x

11x11Before: 11nullshrexschmetricrecon
11 x 11x  11nullshrexpointing
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtract (conditional) x nullshrexschmetricrecon x current schwarzchild metric and its derivatives on the sphere x

11x11After: 11nullshrexschreadclm
11 x 11Before: 11nullshrexpointing
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxPOSTxRECOVERxVARIABLES (conditional) x nullshrexprecalcnsylm x precalculate spherical harmonics in (p,q) coordinates x

11x11Language:11fortran
11 x 11Type: 11function

11

NullSHRExtract (conditional) x nullshrexanaschmetricrecon x current analytic schwarzchild metric and its derivatives on the sphere x

11x11After: 11nullshrexschmetricrecon
11 x 11Before: 11nullshrexpointing
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

NullSHRExtract (conditional) x nullshrexpointing x gives the target variables for the pointers used in extraction x

11x11Before: 11nullshrexextractwt
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtract (conditional) x nullshrexextractwt x computes the extracted variables on the world tube for the current level x

11x11After: 11nullshrexpointing
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x nullshrexwtsphharm x if requested, output spherical harmonics of the quantities computed on the worldtube x

11x11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxANALYSIS (conditional) x wtxwritesphharmwt x if requested, output spherical harmonics of the quantities computed on the nullgrid just off the worldtube x

11x11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxINITIAL (conditional) x nullshrexschprecalc x computes the schwarzchild metric spherical harmonic coefficients x

11x11After: 11nullshrexprecalcnsylm
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullEvolxBoundaryInit (conditional) x nullshrextractinitial x group of the waveform extraction initialization

11x11Storage:11wtxcmpxcoord
11 x 11x  11wtxscalars
11 x 11x  11wtx3vectors
11 x 11x  11wtx4vectors
11 x 11x  11wtxd4vectors
11 x 11x  11wtxdsigmaxvect
11 x 11x  11wtxg1xvect
11 x 11x  11wtxgupxvect
11 x 11x  11wtxjac
11 x 11x  11wtxaffinexmetric
11 x 11x  11wtxdeta0xvect
11 x 11x  11xxwt[2]
11 x 11x  11jxwt[2]

11 x 11x  11jxl

11 x 11x  11betaxwt[2]
11 x 11x  11betaxl
11 x 11x  11uxwt[2]

11 x 11x  11uxl

11 x 11x  11uxxxwt
11 x 11x  11wxwt[2]

11 x 11x  11wxl

11 x 11x  11qxwt[2]
11 x 11Type: 11group

11

NullEvolxBoundaryInit (conditional) x sphericalharmonicreconread x schedule group for reconstructing initial world-tube data x

11x11Before:11nullshrextractinitial
11 x 11Type: 11group

11

NullSHRExtractInitial (conditional) x nullshrexschreadclm x computes the schwarzchild spherical harmonic coefficients x

11x11Before: 11nullshrexschmetricrecon
11 x 11x  11nullshrexmetricreconpast
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial (conditional) x nullshrexschmetricrecon x schwarzchild metric and its derivatives on the sphere x

11x11After: 11nullshrexschreadclm
11 x 11Before: 11nullshrexmetricreconpast
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial (conditional) x nullshrexanaschmetricrecon x analytic schwarzchild metric and its derivatives on the sphere

11x11After: 11nullshrexschmetricrecon
11 x 11Before: 11nullshrexmetricreconpast
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

NullSHRExtractInitial x nullshrexmetricreconpast x copy current metric to past levels x

11x11Before: 11nullshrexpointingxpxp
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

Aliased Functions x

11Alias Name: xxxxxxx11Function Name:
11NullSHRExInitialWT 11NullSHRExInitialWTxp